OpenStudy (joshoyen):

Express each number in terms of i. 1. √-32 2. 2√ -18 3. √ -1/9

jagr2713 (jagr2713):

Hey what do you think we do first?

OpenStudy (joshoyen):

Uh, i really am not sure

jagr2713 (jagr2713):

Well "i" is another term for -1 so we have \[\sqrt{-32}\] what do you think we do with the i?

OpenStudy (joshoyen):

Multiply it with √ -32?

jagr2713 (jagr2713):

Yes so when we do that, it become postive but first, what do we do with the 32?

OpenStudy (joshoyen):

simplify?

jagr2713 (jagr2713):

Yes so what do we get?

OpenStudy (joshoyen):

uhm not sure how to do that, i forgot

jagr2713 (jagr2713):

Well to do that we need to find 2 numbers that multiply to get 32 BUT one of them has to be able to be sqrt.

OpenStudy (joshoyen):

8 and 4

jagr2713 (jagr2713):

Great, can we simplify more?

jagr2713 (jagr2713):

wait no

jagr2713 (jagr2713):

we need one number of the 2 that can be sqrt

jagr2713 (jagr2713):

For example, lets say we have \[\sqrt{18}\] 6,3,9,2 go into it. But only one of them can be sqrt which is 9 \[\sqrt{9}=3\] so we are left with \[3\sqrt{2}\]

jagr2713 (jagr2713):

Do you get it?

OpenStudy (joshoyen):

so it's √ 16 for 32 right?

jagr2713 (jagr2713):

COrrect and sqrt16 is?

OpenStudy (joshoyen):

4

jagr2713 (jagr2713):

so now 4 have been sqrt so we only have 2 left and that stays under the sqrt. And we bring the "i" outside so what would i answer be?

OpenStudy (joshoyen):

Wait what do you mean we have 2 left,

jagr2713 (jagr2713):

we had 16*2 and we only can sqrt 16 so 2 stays

jagr2713 (jagr2713):

16*2=32

OpenStudy (joshoyen):

ohh i see now

OpenStudy (joshoyen):

So it's 4√ 2

jagr2713 (jagr2713):

you forgot the i D:

OpenStudy (joshoyen):

oh lol 4i√ 2

jagr2713 (jagr2713):

Correct :D

jagr2713 (jagr2713):

Ok #2 what would we get?

OpenStudy (joshoyen):

3i√ 2

OpenStudy (joshoyen):

Then would it be 2x3i√ 2

jagr2713 (jagr2713):

Yes so whats 3*2 i

OpenStudy (joshoyen):

6i√ 2

jagr2713 (jagr2713):

Correct :D #3?

OpenStudy (joshoyen):

oh uhm not sure how to do this one

jagr2713 (jagr2713):

what is 9sqrt?

OpenStudy (joshoyen):

3

jagr2713 (jagr2713):

what is -1(-1)

OpenStudy (joshoyen):

1

jagr2713 (jagr2713):

remember -1 is "I" so its i so we have for the answer?

OpenStudy (joshoyen):

1/3i

jagr2713 (jagr2713):

no -1(i) ?

OpenStudy (joshoyen):

oh -1(i) is -1(-1) which is 1?

jagr2713 (jagr2713):

-3(i) = 3i so -1(i) = i

OpenStudy (joshoyen):

ohh okay

jagr2713 (jagr2713):

So \[\frac{ i }{ 3 }\]

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8

OpenStudy (joshoyen):

Thanks so much !

jagr2713 (jagr2713):

If you need more expert help, i would recommend out Qualified helpers :D For only 1.99 you can get expert help and a better understanding of the concept.

OpenStudy (joshoyen):

Oh really? Is it an only pay once thing?

OpenStudy (joshoyen):

or do i have to pay everytime

jagr2713 (jagr2713):

Its a pay once thing, might change over time but you will receive great help and a better understanding

OpenStudy (joshoyen):

Hm I'll look into it, thank you so much!

jagr2713 (jagr2713):

Yes anytime :D

OpenStudy (joshoyen):

Hm I'll look into it, thank you so much!